Kirk Van Houten, who has been married for 23 years, would like to buy his wife an expensive diamond ring with a platinum setting on their 30-year wedding anniversary. Assume that the cost of the ring will be $9,948 in 7 years. Kirk currently has $4,500 to invest. What annual rate of return must Kirk earn on his investment to accumulate enough money to pay for the ring? Click the table icon to view the PVIF table : 


The annual rate of return Kirk must earn on his investment is _______ %. (Round to the nearest integer.)
Answer #1

Here we will use the following formula:

FV = PV * (1 + r%)n

where, FV = Future value = $9948, PV = Present value = $4500, r = rate of interest, n= time period = 7

now, putting theses values in the above equation, we get,

$9948 = $4500 * (1 + r)7

$9948 / $4500 = (1 + r)7

2.21067 = (1 + r)7

(2.21067)1/7 = 1 + r

1.12 = 1 + r

r = 1.12 - 1

r = 0.12 or 12%

So, annual rate of return is 12%